# Design of Experiments for Emulations: A Selective Review from a Modeling Perspective

### v1

@@ Abstract
Space-filling designs are crucial for efficient computer experiments, enabling accurate surrogate modeling and uncertainty quantification in many scientific and engineering applications, such as digital twin systems and cyber-physical systems. In this work, we will provide a comprehensive review on key design methodologies, including Maximin/miniMax designs, Latin hypercubes, and projection-based designs. Moreover, we will connect the space-filling design criteria like the fill distance to Gaussian process performance. Numerical studies are conducted to investigate the practical trade-offs among various design types, with the discussion on emerging challenges in high-dimensional and constrained settings. The paper concludes with future directions in adaptive sampling and machine learning integration, providing guidance for improving computational experiments.
@@ 2 Space-Filling Design
where $x_{ij}$ is the $(i,j)$th entry of $X$ and $u_{ij}$’s are independent random numbers from $[0,1)$. The “lattice sample” due to Patterson (1954) corresponds to $L$ with $u_{ij}=0.5$ for all $(i,j)$’s.
Table 1 displays a Latin hypercube $L$ of $7$ runs for $3$ input variables $X_{1},X_{2},X_{3}$ and a Latin hypercube design $X$ based on $L$. Figure 1 shows the pairwise plot of $X$ and illustrates the one-dimensional uniformity: when the seven points are projected onto each axis, there is exactly one point in each of the seven equally-spaced intervals.
Table 1: A $7\times 3$ Latin hypercube $L$ and a Latin hypercube design $X$ based on $L$
$L$ $X$
@@
Figure 1: The pairwise plot of the Latin hypercube design $X$ in Table 1 for the three input variables $X_{1},X_{2},X_{3}$
@@ 3 Gaussian Process Emulator
It turns out that the predictor $\hat{y}(\bm{x})$, when $\mu=0$, is the same as the regularized kernel regression. For any symmetric positive definite kernel function defined on $\Omega\subset\mathbb{R}^{d}$, $K$ can induce a reproducing kernel Hilbert space, denoted by $\mathcal{H}_{K}$, which satisfies the two conditions: (1) for any $\bm{x}\in\Omega$, $K(\bm{x},\cdot)\in\mathcal{H}_{K}$; and (2) for any $f\in\mathcal{H}_{K}$ and any $\bm{x}\in\Omega$, $\langle f,K(\bm{x},\cdot)\rangle_{\mathcal{H}_{K}}=f(\bm{x})$, also known as the reproducing property. Here, $\langle\cdot,\cdot\rangle_{\mathcal{H}_{K}}$ stands for the inner product of $\mathcal{H}_{K}$ that is deduced from a bilinear operator <,>$\cdot$,$\cdot$>< ⋅ , ⋅ >. For any $f:\Omega\rightarrow\mathbb{R}$ and $f\in\mathcal{H}_{K}$, there exists such a set $\mathcal{X}=\{\bm{x}_{1},\ldots,\bm{x}_{n}\}$ and $\bm{c}\in\mathbb{R}^{n}$ such that $f(\cdot)=\sum_{i=1}^{n}c_{i}K(\bm{x}_{i},\cdot)$. For any $f,g\in\mathcal{H}_{K}$, which can be written into $f,g\in\mathcal{F}$ with $f=\sum_{i=1}^{n}c_{i}K(\bm{x}_{i},\cdot)$ and $g=\sum_{j=1}^{m}d_{j}K(\bm{z}_{j},\cdot)$, the inner product and the norm based on it are defined by
$\langle f,g\rangle_{\mathcal{H}_{K}}=\sum_{i=1}^{n}\sum_{j=i}^{m}c_{i}d_{j}K(\bm{x}_{i},\bm{z}_{j}),\quad||f||_{\mathcal{H}_{K}}=\sum_{i,j=1}^{n}c_{i}c_{j}K(\bm{x}_{i},\bm{x}_{j})=\bm{c}^{\top}\bm{K}\bm{c}.$
@@ 4.1 Fill and Separation Distance
$\|D^{\bm{\alpha}}f-D^{\bm{\alpha}}\hat{y}\|_{L_{\infty}(\Omega)}\leq Ch_{\mathcal{X},\Omega}^{(k+\nu)/2-\|\bm{\alpha}\|_{1}}||f||_{\mathcal{H}_{K}}.$
Remark: There are several technical definitions that might be less familar to readers. The notation $\bm{\alpha}$ is derivative order for the $d$-dimensions and $D^{\bm{\alpha}}$ denotes the derivative operator of the orders and dimensions specified by $\bm{\alpha}$. A conditionally positive definite kernel is a function that, when used to create the Gram matrix (such as the kernel matrix in kriging), results in a matrix that is conditionally positive definite, meaning it is positive semi-definite when the sum of the coefficients is zero. The order $m$ refers to the highest degree of polynomials that can be reproduced by the kernel associated with the conditionally positive definite kernel. Thus, conditionally positive definite kernels include all positive definite kernels, such as the Gaussian and Matérn kernels we have showed earlier. The set $C_{\nu}^{k}(\mathbb{R}^{d})$ is Hölder space consisting of all functions $f\in C^{k}(\mathbb{R}^{d})$ (continuous to the $k$th degree) whose derivatives of order $k$ satisfy $D^{\bm{\alpha}}f(\bm{x})=O(\|\bm{x}\|_{2}^{\nu})$ for $\|\bm{x}\|_{2}^{\nu}\to 0$. If $\Omega$ is bounded and satisfies an interior cone condition, simply put, it implies if $\bm{x}\in\Omega$, a cone with $\bm{x}$ as vertex is also in $\Omega$. Here $C$ and $h_{0}$ are both some constant that only depend on $\Omega$. Theorem 1 is general and covers the error bound of the derivatives with order $\|\bm{\alpha}\|_{1}\leq k/2$ of $f-\hat{y}$ for a wide range of kernel functions. As a special case,
$||f-\hat{y}||_{L_{\infty}(\Omega)}\leq Ch_{\mathcal{X},\Omega}^{(k+\nu)/2}||f||_{\mathcal{H}_{K}}.$
@@
$\max_{\mathcal{X}\subset\Omega}\min_{\bm{x},\bm{x}^{\prime}\in\mathcal{X}}||\bm{x}-\bm{x}^{\prime}||_{2},$
@@
where $\text{dist}(\bm{z},\mathcal{X})=\min_{\bm{x}\in\mathcal{X}}||\bm{z}-\bm{x}||_{2}$. Pronzato (2017) further explained that
$\min_{\mathcal{X}\subset\Omega}\max_{\bm{z}\in\Omega}\text{dist}(\bm{z},\mathcal{X})=\max_{\bm{x}\in\Omega}\min_{i=1,\ldots,n}||\bm{x}-\bm{x}_{i}||_{2},$
which is a direct approximation of the fill distance in (4).
@@ 5.1 Distance-based Latin Hypercube Designs
where $q$ is an integer with $q=1$ and $q=2$ corresponding to the rectangular and Euclidean distances, respectively. Note that the distance in (5) is just a commonly used distance for the continuous variables. One can also consider other types of distance. Let ${\chi}$ be the design space throughout.
@@
$\max_{\mbox{\boldmath{$x$}}\in\mbox{ \boldmath${\chi}$}}d_{q}(\mbox{\boldmath{$x$}},\mbox{\boldmath{$X$}}),$
where $d_{q}(\mbox{\boldmath{$x$}},\mbox{\boldmath{$X$}})$ denotes the $L_{q}$ distance between $x$ and its nearest neighbor in $X$, defined as
@@
Figure 2: The performance in terms of (7) and (6) of maximin Latin hypercube designs generated by the $maximinSLHD,maximinESE\_LHS,maximinSA\_LHS$ functions
@@ 5.2 Orthogonal and Nearly Orthogonal Latin Hypercube Designs
Extensive research has been conducted on the construction of orthogonal and nearly orthogonal Latin hypercube designs. The foundation for this research was established by Ye (1998), who pioneered the construction of orthogonal Latin hypercubes with $n=2^{m}$ or $12^{m}+$ 1 runs and $k=2m-2$ factors, where $m\geq 2$. Subsequent work has expanded upon these initial results in several directions: Cioppa and Lucas (2007) extended Ye’s construction to accommodate more columns for given run sizes; Steinberg and Lin (2006) developed orthogonal Latin hypercube designs with run sizes $n=2^{2^{m}}$ by rotating groups of factors in two-level regular fractional factorial designs; Pang et al. (2009) generalized this approach to $p^{2^{m}}$ runs (where $p$ is prime) with up to $(p^{2^{m}}-1)/(p-1)$ factors using $p$-level regular factorial designs; Lin (2008) introduced an algorithmic approach for small run sizes ($n\leq 20$) by sequentially adding columns to existing designs. Recent advances have focused on creating orthogonal Latin hypercube designs with more flexible run sizes and higher factor-to-run-size ratios. Of particular note are the general and computationally efficient methods proposed by Lin et al. (2010) and Sun and Tang (2017a), which encompass the results presented in Table 3. Additional construction methods can be found in the works of Lin et al. (2009), Sun et al. (2009), Georgiou (2009), Sun et al. (2010), Yang and Liu (2012), and Sun and Tang (2017b), demonstrating the continued development and refinement of orthogonal Latin hypercube design construction methods.
@@
Table 3: The best lower bound $k$ on the maximum number $k^{*}$ of factors in ${\mbox{OLH}}(n,k^{*})$ for $24512\geq n>$ 24
@@ 5.3 Projection-based Latin Hypercube Designs
Let $S$ be a set of $s$ levels which are taken to be $1,2,\ldots,s$ in this article. An $s$-level orthogonal array of $n$ runs, $d$ factors and strength $t$, denoted by ${\mbox{OA}}(n,s^{d},t)$ is an $n\times d$ array $D$ with entries from $S$ with the property that every $n\times t$ subarray of $D$ contains all possible level combinations of $t$ factors exactly $\lambda$ times as a row. By the definition of orthogonal arrays, a Latin hypercube of $n$ runs for $d$ factors is an ${\mbox{OA}}(n,n^{d},1)$.
The construction of orthogonal array-based Latin hypercubes, as proposed in Tang (1993), proceeds as follows. Let $D$ be an ${\mbox{OA}}(n,s^{d},t)$. For each column of $D$ and $m=1,\ldots,s$, replace the $n/s$ positions containing the entry $m$ by a random permutation of $(m-1)n/s+1,(m-1)n/s+2,\ldots,mn/s$. Denote the resulting design after this replacement procedure as $L$. A corresponding orthogonal array-based Latin hypercube design in the space $[0,1)^{d}$ can then be obtained via (1). In addition to ensuring one-dimensional uniformity, an ${\mbox{OA}}(n,s^{d},t)$-based Latin hypercube possesses the $t$-dimensional projection property. Specifically, when projected onto any $t$ columns, the design places exactly $\lambda=n/s^{t}$ points in each of the $s^{t}$ cells $\mathcal{P}^{t}$ where $\mathcal{P}=\{[0,1/s],[1/s,2/s),\ldots,[1-1/s,1)\}$. Example 5.2 illustrates the two-dimensional projection property of an ${\mbox{OA}}(9,3^{4},2)$-based Latin hypercube. To enhance projection properties, He and Tang (2013) introduced the concept of strong orthogonal arrays, which serve as a foundation for constructing Latin hypercube designs with superior space-filling characteristics. This has spurred active research on the construction of strong orthogonal arrays; see Chen and Tang (2024) and references therein for further developments.
@@
Table 4 presents an orthogonal array-based Latin hypercube, $L$, constructed from the orthogonal array ${\mbox{OA}}(9,3^{4},2)$. Figure 3 illustrates the pairwise scatter plot of this Latin hypercube. Each subplot contains exactly one point within each of the nine dot-dash line boxes, demonstrating the design’s uniformity in two-dimensional projections.
@@ 5.4.1 Designs with Many Levels
[…]
@@ 5.4.2 Low-discrepancy Sequence
[…]
$|\mu-\hat{\mu}|\leq V(f)D^{*}(\mbox{ \boldmath${\chi}$}),$
where $V(f)$ is the variation of $f$ in the sense of Hardy and Krause and $D^{*}(\mbox{ \boldmath${\chi}$})$ is the star discrepancy,
$D^{*}(\mbox{ \boldmath${\chi}$})=\max_{\mbox{\boldmath{$x$}}\in\mbox{ \boldmath${\chi}$}}\left|\frac{N(\mbox{ \boldmath${\chi}$},J_{\mathbf{x}})}{n}-\hbox{Vol}({J_{\mathbf{x}}})\right|,$
with $J_{\mathbf{x}}=[0,{\mathbf{x}})$ being the interval of $[0,x_{1})\times[0,x_{2})\times\cdots\times[0,x_{d})$, $N(\mbox{ \boldmath${\chi}$},J_{\mathbf{x}})$ being the number of points of ${\chi}$ falling in $J_{\mathbf{x}}$, and $\hbox{Vol}(J_{\mathbf{x}})$ being the volume if the interval $J_{\mathbf{x}}$. The star discrepancy measures how well the points in ${\chi}$ approximate the uniform distribution by comparing the proportion of points in subregions of ${\chi}$ to their theoretical volumes. Quasi-Monte Carlo methods seek to generate points that minimize this discrepancy. A sequence $S$ is considered a low-discrepancy sequence if its first $n$ points satisfy
$D^{*}(\mbox{ \boldmath${\chi}$})=O(n^{-1}\hbox{log}(n)^{d}),$
@@ Acknowledgment
This section presents four test function based simulators used in Sections 4 for performance comparison of the competing design criteria for global fitting respectively. All test functions were scaled such that the inputs are in $[0,1]^{d}$.

### v2

@@ Abstract
Space-filling designs are crucial for efficient computer experiments, enabling accurate surrogate modeling and uncertainty quantification in many scientific and engineering applications, such as digital twin systems and cyber-physical systems. In this work, we will provide a comprehensive review of key design methodologies, including Maximin/miniMax designs, Latin hypercubes, and projection-based designs. Moreover, we will connect the space-filling design criteria, like the fill distance, to the Gaussian process performance. Numerical studies are conducted to investigate the practical trade-offs among various design types, with a discussion on emerging challenges in high-dimensional and constrained settings. The paper concludes with future directions in adaptive sampling and machine learning integration, providing guidance for improving computational experiments.
@@ 2 Space-Filling Design
[…]
@@
[…]
@@ 3 Gaussian Process Emulator
It turns out that the predictor $\hat{y}(\bm{x})$, when $\mu=0$, is the same as the regularized kernel regression. For any symmetric positive definite kernel function defined on $\Omega\subset\mathbb{R}^{d}$, $K$ can induce a reproducing kernel Hilbert space, denoted by $\mathcal{H}_{K}$, which satisfies the two conditions: (1) for any $\bm{x}\in\Omega$, $K(\bm{x},\cdot)\in\mathcal{H}_{K}$; and (2) for any $f\in\mathcal{H}_{K}$ and any $\bm{x}\in\Omega$, $\langle f,K(\bm{x},\cdot)\rangle_{\mathcal{H}_{K}}=f(\bm{x})$, also known as the reproducing property. Here, $\langle\cdot,\cdot\rangle_{\mathcal{H}_{K}}$ stands for the inner product of $\mathcal{H}_{K}$ that is deduced from a bilinear operator $\cdot$$\cdot$>. For any $f:\Omega\rightarrow\mathbb{R}$ and $f\in\mathcal{H}_{K}$, there exists such a set $\mathcal{X}=\{\bm{x}_{1},\ldots,\bm{x}_{n}\}$ and $\bm{c}\in\mathbb{R}^{n}$ such that $f(\cdot)=\sum_{i=1}^{n}c_{i}K(\bm{x}_{i},\cdot)$. For any $f,g\in\mathcal{H}_{K}$, which can be written into $f,g\in\mathcal{F}$ with $f=\sum_{i=1}^{n}c_{i}K(\bm{x}_{i},\cdot)$ and $g=\sum_{j=1}^{m}d_{j}K(\bm{z}_{j},\cdot)$, the inner product and the norm based on it are defined by
[…]
@@ 4.1 Fill and Separation Distance
[…]
Remark: There are several technical definitions that might be less familiar to readers. The notation $\bm{\alpha}$ is derivative order for the $d$-dimensions and $D^{\bm{\alpha}}$ denotes the derivative operator of the orders and dimensions specified by $\bm{\alpha}$. A conditionally positive definite kernel is a function that, when used to create the Gram matrix (such as the kernel matrix in kriging), results in a matrix that is conditionally positive definite, meaning it is positive semi-definite when the sum of the coefficients is zero. The order $m$ refers to the highest degree of polynomials that can be reproduced by the kernel associated with the conditionally positive definite kernel. Thus, conditionally positive definite kernels include all positive definite kernels, such as the Gaussian and Matérn kernels we have showed earlier. The set $C_{\nu}^{k}(\mathbb{R}^{d})$ is Hölder space consisting of all functions $f\in C^{k}(\mathbb{R}^{d})$ (continuous to the $k$th degree) whose derivatives of order $k$ satisfy $D^{\bm{\alpha}}f(\bm{x})=O(\|\bm{x}\|_{2}^{\nu})$ for $\|\bm{x}\|_{2}^{\nu}\to 0$. If $\Omega$ is bounded and satisfies an interior cone condition, simply put, it implies if $\bm{x}\in\Omega$, a cone with $\bm{x}$ as vertex is also in $\Omega$. Here $C$ and $h_{0}$ are both some constant that only depend on $\Omega$. Theorem 1 is general and covers the error bound of the derivatives with order $\|\bm{\alpha}\|_{1}\leq k/2$ of $f-\hat{y}$ for a wide range of kernel functions. As a special case,
[…]
@@
[…]
@@
[…]
which is a direct approximation of the fill distance in (4). Therefore, miniMax design aims to minimize the fill distance.
@@ 5.1 Distance-based Latin Hypercube Designs
where $q$ is an integer with $q=1$ and $q=2$ corresponding to the rectangular and Euclidean distances, respectively. Note that the distance in (5) is just a commonly used distance for the continuous variables. One can also consider other types of distance. Let $\mathcal{X}$ be the design space throughout.
@@
$\max_{\mbox{\boldmath{$x$}}\in\mathcal{X}}d_{q}(\mbox{\boldmath{$x$}},\mbox{\boldmath{$X$}}),$
[…]
@@
Figure 2: The performance in terms of (7) and (6) of maximin Latin hypercube designs generated by the maximinSLHD,maximinESE_LHS,maximinSA_LHSmaximinSLHD,maximinESE\_LHS,maximinSA\_LHS functions
@@ 5.2 Orthogonal and Nearly Orthogonal Latin Hypercube Designs
Extensive research has been conducted on the construction of orthogonal and nearly orthogonal Latin hypercube designs. The foundation for this research was established by Ye (1998), who pioneered the construction of orthogonal Latin hypercubes with $n=2^{m}$ or $2^{m}+1$ runs and k=2m2k=2m-2 factors, where $m\geq 2$. Subsequent work has expanded upon these initial results in several directions: Cioppa and Lucas (2007) extended Ye’s construction to accommodate more columns for given run sizes; Steinberg and Lin (2006) developed orthogonal Latin hypercube designs with run sizes $n=2^{2^{m}}$ by rotating groups of factors in two-level regular fractional factorial designs; Pang et al. (2009) generalized this approach to $p^{2^{m}}$ runs (where $p$ is prime) with up to $(p^{2^{m}}-1)/(p-1)$ factors using $p$-level regular factorial designs; Lin (2008) introduced an algorithmic approach for small run sizes ($n\leq 20$) by sequentially adding columns to existing designs. Recent advances have focused on creating orthogonal Latin hypercube designs with more flexible run sizes and higher factor-to-run-size ratios. Of particular note are the general and computationally efficient methods proposed by Lin et al. (2010) and Sun and Tang (2017a), which encompass the results presented in Table 3. Additional construction methods can be found in the works of Lin et al. (2009), Sun et al. (2009), Georgiou (2009), Sun et al. (2010), Yang and Liu (2012), and Sun and Tang (2017b), demonstrating the continued development and refinement of orthogonal Latin hypercube design construction methods.
@@
Table 3: The best lower bound $k$ on the maximum number $k^{*}$ of factors in ${\mbox{OLH}}(n,k^{*})$ for $512\geq n>24$
@@ 5.3 Projection-based Latin Hypercube Designs
[…]
@@
[…]
@@ 5.4.1 Designs with Many Levels
Bingham et al. (2009) argued that in many practical applications, it is not necessary for the number of levels to equal the run size. Instead, they proposed the use of orthogonal and nearly orthogonal designs for computer experiments, which allow factors to have repeated levels while maintaining zero or low correlations. This approach encompasses a broad class of orthogonal designs, including two-level orthogonal designs and orthogonal Latin hypercubes as special cases. For construction methods of orthogonal and nearly orthogonal designs, see Bingham et al. (2009) and Georgiou et al. (2014). Table 5 presents a $4\times 2$ orthogonal design $\mbox{\boldmath{$D$}}_{1}$ and a $8\times 4$ orthogonal design $\mbox{\boldmath{$D$}}_{2}$ where in $\mbox{\boldmath{$D$}}_{1}$, $x_{1}$ and $x_{2}$ is a permutation of $\{1,2\}$, and in $\mbox{\boldmath{$D$}}_{2}$, $x_{1},x_{2},x_{3},x_{4}$ is a permutation of $\{1,2,3,4\}$.
@@ 5.4.2 Low-discrepancy Sequence
Another widely used class of space-filling designs is quasi-random low-discrepancy sequences, which measure uniformity based on discrepancy, or the deviation from a perfectly uniform distribution. These sequences originate from numerical integration problems, where the goal is to estimate the expectation $\mu$ of $y=f(\mbox{\boldmath{$x$}})$ with $f$ being known and $\mbox{\boldmath{$x$}}=(x_{1},\ldots,x_{d})$ having a uniform distribution in the design space $\mathcal{X}=[0,1)^{d}$ and $y\in R$. Given a set of $n$ points in the design space $\mathcal{X}$, and using $\hat{\mu}$ to approximate $\mu$, the integration error bound is determined by the Koksma-Hlawka inequality,
$|\mu-\hat{\mu}|\leq V(f)D^{*}(\mathcal{X}),$
where $V(f)$ is the variation of $f$ in the sense of Hardy and Krause and $D^{*}(\mathcal{X})$ is the star discrepancy,
$D^{*}(\mathcal{X})=\max_{\mbox{\boldmath{$x$}}\in\mathcal{X}}\left|\frac{N(\mathcal{X},J_{\mathbf{x}})}{n}-\hbox{Vol}({J_{\mathbf{x}}})\right|,$
with $J_{\mathbf{x}}=[0,{\mathbf{x}})$ being the interval of $0,x_{1})\times[0,x_{2})\times\cdots\times[0,x_{d})$, $N(\mathcal{X},J_{\mathbf{x}})$ being the number of points of $\mathcal{X}$ falling in $J_{\mathbf{x}}$, and $\hbox{Vol}(J_{\mathbf{x}})$ being the volume if the interval $J_{\mathbf{x}}$. The star discrepancy measures how well the points in $\mathcal{X}$ approximate the uniform distribution by comparing the proportion of points in subregions of $\mathcal{X}$ to their theoretical volumes. Quasi-Monte Carlo methods seek to generate points that minimize this discrepancy. A sequence $S$ is considered a low-discrepancy sequence if its first $n$ points satisfy
$D^{*}(\mathcal{X})=O(n^{-1}\hbox{log}(n)^{d}),$
@@ Acknowledgment
This section presents four test function based simulators used in Sections 4 for performance comparison of the competing design criteria for global fitting respectively. All test functions were scaled such that the inputs are in $0,1]^{d}$.